OpenStudy (anonymous):

Which expression is equivalent to (2x - 5) - ( 3x - 8) A: -x - 13 B: -x + 3 C: 2x - 16 D: 5x - 3

OpenStudy (anonymous):

(2x - 5) - ( 3x - 8) = 3-x

OpenStudy (anonymous):

thats 2x - 5 - 3x + 8 = -x + 3 or 3-x

OpenStudy (anonymous):

Okay. Thanks(: